OpenStudy (anonymous):

how do you solve this equation√8k=k

OpenStudy (anonymous):

sqrt(8k) = k 8k = k^2 0 = k^2 - 8k 0 = k(k-8) k = 0 or k = 8

OpenStudy (anonymous):

we square everything. put it to one side. pull k out as a common factor.

OpenStudy (anonymous):

thanks how do you solve this equation √3-2x=√1-3x

OpenStudy (anonymous):

is the squareroot over the whole thing. use the, \sqrt{3-2x} in the equation thing if you mean that

OpenStudy (anonymous):

yes the square root is over the whole thing

OpenStudy (anonymous):

square both sides then.